# Say More with Less: Variable-Frame-Rate Speech Tokenization via Adaptive Clustering and Implicit Duration Coding

### v1

@@ 4.4 TTS Language Modeling
To evaluate the compatibility of VARSTok with downstream speech LMs, we trained an TTS model based on a decoder-only autoregressive architecture. Similar to VALL-E (Chen et al. 2025b), given an input text and a 3-second speech prompt from the target speaker, the model autoregressively generates speech tokens, which are then directly decoded into waveform using the the decoder of the pretrained speech tokenizer. As shown in Table 3, VARSTok with $\tau=0.8$ and $S_{\max}=4$ achieves the best overall performance despite using fewer tokens on average, improving naturalness while reducing WER compared to the 40 Hz WavTokenizer baseline. Even at a lower token rate of 30.95 Hz, our model achieves comparable WER and higher UTMOS, demonstrating that VARSTok enhances both efficiency and generation quality. We note a slight decrease in the objective speaker similarity score as the token rate drops. However, such metrics are known to sometimes diverge from human perception. Subjective results from human evaluators confirm that VARSTok matches the baseline in perceived speaker similarity and naturalness, indicating the minor drop in the objective SIM score is not perceptually significant. These results reinforces the overall superiority of our dynamic tokenization approach for enabling token-efficient and high-fidelity speech synthesis. Reprensentative speech samples are provided in the Appendix A.
@@ 5 Conclusion
A. Speech Samples
Speech samples are available at our demo page https://anonymous.4open.science/w/VARSTok-demo-8E60/ .

### v2

@@ 4.4 TTS Language Modeling
To evaluate the compatibility of VARSTok with downstream speech LMs, we trained an TTS model based on a decoder-only autoregressive architecture. Similar to VALL-E (chen2025neural), given an input text and a 3-second speech prompt from the target speaker, the model autoregressively generates speech tokens, which are then directly decoded into waveform using the the decoder of the pretrained speech tokenizer. As shown in Table 3, VARSTok with $\tau=0.8$ and $S_{\max}=4$ achieves the best overall performance despite using fewer tokens on average, improving naturalness while reducing WER compared to the 40 Hz WavTokenizer baseline. Even at a lower token rate of 30.95 Hz, our model achieves comparable WER and higher UTMOS, demonstrating that VARSTok enhances both efficiency and generation quality. We note a slight decrease in the objective speaker similarity score as the token rate drops. However, such metrics are known to sometimes diverge from human perception. Subjective results from human evaluators confirm that VARSTok matches the baseline in perceived speaker similarity and naturalness, indicating the minor drop in the objective SIM score is not perceptually significant. These results reinforces the overall superiority of our dynamic tokenization approach for enabling token-efficient and high-fidelity speech synthesis. Reprensentative speech samples are provided in the Appendix A.
@@ 5 Conclusion
6 Acknowledgments
This work was partially funded by the National Nature Science Foundation of China under Grant 62301521 and the Anhui Province Major Science and Technology Research Project under Grant S2023Z20004.
Technical Appendix
A. Speech Samples and Implementation Codes
Speech samples and implementation codes are available at https://anonymous.4open.science/w/VARSTok-demo-8E60/ .